OpenStudy (anonymous):

Solve. 7^x – 5 = 7^4x + 7

OpenStudy (anonymous):

since base is same i.e. 7 and LHS = RHS => x-5 = 4x +7 => -5 - 7 = 4x - x => -12 = 3x => -12/3 = x => -4 = x

OpenStudy (anonymous):

explained nicely....

OpenStudy (anonymous):

Thanks....☺